OpenStudy (fluttershyk):

(I have a chart) Drink Survey Drink Number of shoppers Who preferred it A 10 B 15 C 7 D 3 E 6 What is the probability that 1 shopper, selected at random, preferred neither Drink E nor Drink C? 35/41 13/41 28/41 13/28

OpenStudy (fluttershyk):

@Cas_fangirl_14 @rebeccaxhawaii @ILovePuppiesLol

rebeccaxhawaii (rebeccaxhawaii):

10+15+7+3+6=?

OpenStudy (fluttershyk):

41

rebeccaxhawaii (rebeccaxhawaii):

okay so drink c is 7 and e is 6 so 7+6=13 what is 41-13=?

OpenStudy (fluttershyk):

28/41

rebeccaxhawaii (rebeccaxhawaii):

correct c: